# UniHGKR: Unified Instruction-aware Heterogeneous Knowledge Retrievers

### v1

@@ 6.3 Extending UniHGKR to LLM Retrievers
Recent works, such as E5-mistral-7B Wang et al. (2024) and LLARA Li et al. (2023a), have explored converting decoder-only LLMs into dense retrievers, leveraging their extensive pre-trained knowledge to achieve improvements on various IR tasks. Our UniHGKR framework is plug-and-play and can seamlessly adapt to training LLM retrievers by adjusting the training objectives. To demonstrate this, we adapt LLARA’s methods to train our UniHGKR-7B retrievers based on the CompMix-IR. More adaptation details are in Appendix F.
Table 5 presents the evaluation results of UniHGKR-7B alongside other LLM-based baselines, including E5-mistral-7B, LLARA-passage (LLARA-pretrain fine-tuned on MS MARCO passage), and LLARA-finetuned (LLARA-pretrain fine-tuned on CompMix-IR). Although UniHGKR-7B performs slightly lower than LLARA-passage on the Table-Hit metric, it significantly outperforms the baselines and UniHGKR-base on the other seven metrics, achieving a 23.91% relative improvement in MRR@100 and 18.08% in Hit@5. These results further validate the effectiveness and scalability of our UniHGKR.
@@ 6.4 Employing UniHGKR on QA systems
In this section, we explore the application of UniHGKR retrievers in open-domain QA systems over heterogeneous sources. We select a popular task, ConvMix Christmann et al. (2022b), which is a conversational format variant of the CompMix. This task is more challenging because it requires systems to consider both the current turn’s question and the dialogue history. Baseline models such as QuReTeC Voskarides et al. (2020), CONVINSE Christmann et al. (2022b), and EXPLAIGNN Christmann et al. (2023), along with their results, are sourced from the ConvMix leaderboard [^4] https://convinse.mpi-inf.mpg.de/. We evaluate our systems using the same metrics as them: P@1 (Precision at 1) and MRR.
As shown in Table 6, by replacing retrievers with our UniHGKR models in baseline systems, we observe significant improvements in QA performance. Specifically, compared to CONVINSE, which uses the same reader (FiD, Fusion-in-Decoder Izacard and Grave (2021)) as we do, using UniHGKR-base as the retriever achieves an absolute improvement of up to 8.80 points in MRR, while UniHGKR-7B achieves an improvement of up to 13.60 points in MRR. Compared to the current SOTA system, EXPLAIGNN, which uses a graph neural network (GNN) as a reader, our system surpasses it by up to 4.80 points in MRR ans 4.10 points in P@1, setting a new SOTA performance for the ConvMix task. These results further validate the effectiveness of UniHGKR and also indicate that the retrieval component is a significant factor limiting the performance of current open-domain QA systems on heterogeneous data.
Methods Retriever Reader P@1 MRR QuReTeC QuReTeC FiD 28.2 28.9 CONVINSE CLOCQ+BM25 FiD 34.3 37.8 EXPLAIGNN CLOCQ+BM25 GNN 40.6 47.1 Ours UniHGKR-base FiD 42.4 46.6 Abs. gain +8.10 +8.80 UniHGKR-7B FiD 46.5 51.4 Abs. gain +12.20 +13.60 SOTA gain +4.10 +4.80
Table 6: The performance of the QA system employing the UniHGKR retrievers on the ConvMix dataset.
@@ 8 Limitations
In our study, the CompMix-IR dataset is primarily sourced from Wikidata knowledge graphs and Wikipedia, including infoboxes, tables, and text, but it is limited to five domains: books, movies, music, television series, and football. This may restrict the model’s generalization capabilities. Additionally, while UniHGKR incorporates diverse user instructions, it does not cover all scenarios in heterogeneous information retrieval. For instance, users might want to instruct the retriever to return a combination of evidence from multiple knowledge sources, such as text and tables, or a mix of KG triples, tables, and text, as noted in Christmann et al. (2022b). Exploring these user-defined combinations remains an area for future work. We will open-source our instruction set, CompMix-IR corpus, and UniHGKR model and code, encouraging the community to contribute more retrieval tasks with human-written instructions to assess whether broader instruction coverage enhances performance
@@ Appendix A CompMix-IR Example
We present some question-answer examples from the CompMix-IR dataset in Table 7, while Table  8 shows the statistics of the CompMix-IR QA set. Table 9 provides examples of evidence for the four types of knowledge.
@@ 19 August 1988
Table 9: Evidence examples from the CompMix-IR dataset.
@@ Appendix E Retrieving Robustness of UniHGKR
Appendix F Detailed Description of UniHGKR-7B Adaptation
@@
In our UniHGKR-7B training, we initialize the model weights from the LLaRA-pretrain. Our input sequence $S$ is the linearized structured data $d_{i}$. We adapt EBAE to reconstruct $d_{i}$ and EBAR to predict the corresponding natural language sentence $t_{i}$. Here $\langle d_{i},t_{i}\rangle$ from Data-Text Pairs $\mathcal{D}$. This process essentially implements Stages 1 and 2 of our UniHGKR training framework. For task fine-tuning (Stage 3), we use the same training methods as the UniHGKR-base models (BERT-based), including the instruction set and positive/negative sampling strategies (see Section 4.2).

### v2

@@ 6.3 Extending UniHGKR to LLM Retrievers
Recent works, such as E5-mistral-7B Wang et al. (2024) and LLARA Li et al. (2023a), have explored converting decoder-only LLMs into dense retrievers, leveraging their extensive pre-trained knowledge to achieve improvements on various IR tasks. Our UniHGKR framework is plug-and-play and can seamlessly adapt to training LLM retrievers by adjusting the training objectives. To demonstrate this, we adapt UniHGKR framework to train our UniHGKR-7B retrievers based on the LLARA architecture. More adaptation details are in Appendix G.
Table 5 presents the evaluation results of UniHGKR-7B alongside other LLM-based baselines, including E5-mistral-7B, LLARA-passage (LLARA-pretrain fine-tuned on MS MARCO passage), and LLARA-finetuned (LLARA-pretrain fine-tuned on CompMix-IR). We can observe that our UniHGKR-7B significantly outperforms the LLM-based baselines and UniHGKR-base, achieving SOTA performance on all metrics across two scenarios. In particular, it achieves a 23.91% relative improvement on the MRR@100 metric in Scenario 1 and reaches 49.57 on the Table-Hit in Scenario 2. These results further validate the effectiveness and scalability of our UniHGKR method, as well as the potential of LLMs as retrievers.
@@ 6.4 Employing UniHGKR on QA systems
In this section, we explore the application of UniHGKR retrievers in open-domain QA systems over heterogeneous sources. We select a popular task, ConvMix Christmann et al. (2022b), which is a conversational format variant of the CompMix. This task is more challenging because it requires systems to consider both the current turn’s question and the dialogue history. Baseline models such as QuReTeC Voskarides et al. (2020), CONVINSE Christmann et al. (2022b), and EXPLAIGNN Christmann et al. (2023), along with their results, are sourced from the ConvMix leaderboard [^4] https://convinse.mpi-inf.mpg.de/. Note that in the QA system experiment, we replace the entire retrieval component (e.g., CLOCQ+BM25) of the baseline with our UniHGKR model, not just BM25. The retrieval component of EXPLAIGNN and CONVINSE can be seen as a combination of coarse retrieval (CLOCQ) and re-ranking (BM25). All baseline methods and our UniHGKR use the same corpus to ensure a fair comparison. In the reasoning part after retrieval, we follow CONVINSE and use Fusion-in-Decoder (FiD) Izacard and Grave (2021) as the reader. We input the top 100 relevant evidences returned by the retriever into the reader for inference. Then we evaluate the output of the reader as the performance of the QA system using the same metrics as baselines: P@1 (Precision at 1) and MRR.
As shown in Table 6, by replacing retrievers with our UniHGKR models in baseline systems, we observe significant improvements in QA performance. Specifically, compared to CONVINSE, which uses the same reader FiD as we do, using UniHGKR-base as the retriever achieves an absolute improvement of up to 8.80 points in MRR, while UniHGKR-7B achieves an improvement of up to 13.60 points in MRR. Compared to the current SOTA system, EXPLAIGNN, which uses a graph neural network (GNN) as a reader, our system surpasses it by up to 4.30 points in MRR and 5.90 points in P@1, setting a new SOTA performance for the ConvMix dataset. These results further validate the effectiveness of UniHGKR and also indicate that the retrieval component is a significant factor limiting the performance of current open-domain QA systems on heterogeneous data.
Methods Retriever Reader P@1 MRR BM25+FiD BM25 FiD 25.3 27.5 QuReTeC QuReTeC FiD 28.2 28.9 CONVINSE CLOCQ+BM25 FiD 34.3 37.8 EXPLAIGNN CLOCQ+BM25 GNN 40.6 47.1 Ours UniHGKR-base FiD 42.4 46.6 Abs. gain +8.10 +8.80 UniHGKR-7B FiD 46.5 51.4 Abs. gain +12.20 +13.60 SOTA gain +5.90 +4.30
Table 6: The QA performance of systems using the UniHGKR retriever and baselines on the ConvMix dataset. ‘Abs. gain’ represents the absolute improvement brought by the retriever under the same Reader setting (compared to CONVINSE). ‘SOTA gain’ indicates the absolute improvement over the previous SOTA system.
@@ 8 Limitations
In our study, the CompMix-IR dataset is primarily sourced from Wikidata knowledge graphs and Wikipedia, including infoboxes, tables, and text, but it is limited to five domains: books, movies, music, television series, and football. This may restrict the model’s generalization capabilities. Additionally, while UniHGKR incorporates diverse user instructions, it does not cover all scenarios in heterogeneous information retrieval. For instance, users might want to instruct the retriever to return a combination of evidence from multiple knowledge sources, such as text and tables, or a mix of KG triples, tables, and text, as noted in Christmann et al. (2022b). Exploring these user-defined combinations remains an area for future work. In addition, more modalities such as image, audio and interleaved image and text Xu et al. (2024b) can be considered and incorporated in the retrieving process of UniHGKR in future. We will open-source our instruction set, CompMix-IR corpus, and UniHGKR model and code, encouraging the community to contribute more retrieval tasks with large-scale human-written instructions Xu et al. (2024a) to assess whether broader instruction coverage enhances performance.
@@ Appendix A CompMix-IR Example
A.1 Heterogeneous Evidence Examples
Table 7 provides linearized heterogeneous examples of evidence for the four types of knowledge. Table 8 provides examples of evidence with full annotation information.
Table 7: Evidence examples from the CompMix-IR corpus.
Text evidence: { "linearized evidence text": "Museum of Modern Art, Its first successful loan exhibition was in November 1929, displaying paintings by Van Gogh, Gauguin, Cézanne, and Seurat.", "wikidata entities": [ { "id": "Q34013", "label": "Georges Seurat" }, { "id": "Q17437796", "label": "featured article" }, …], "disambiguations": [ [ "1929", "1929-01-01T00:00:00Z" ], [ "painting", "Q11629" ], [ "Van Gogh", "Q17437796" ], … ], "retrieved for entity": { "id": "Q188740", "label": "Museum of Modern Art" }, "source": "text" } KG evidence: {"linearized evidence text": "Transformers, genre, action film", "wikidata entities": [ { "id": "Q171453", "label": "Transformers" }, { "id": "Q188473", "label": "action film" } ], "disambiguations": [ [ "Transformers", "Q171453" ],…], "source": "kg"} Table evidence:
{ "linearized evidence text": "Surrey Scorchers, Season is 2008–09, Division is BBL, Tier is I, Regular Season is 4th, Post-Season is 33, Trophy is 21, Cup is 12, Head Coach is 42",
"wikidata entities": [ { "id": "Q3645013", "label": "2008–09 British Basketball League season" }, { "id": "Q269597", "label": "Surrey Scorchers" }, { "id": "Q23276", "label": "Surrey"}, … ], "disambiguations": [ [ "2008", "2008-01-01T00:00:00Z" ], [ "2008–09", "Q3645013" ], …], "retrieved for entity": { "id": "Q269597" }, "source": "table"} Infobox evidence: { "linearized evidence text": "Older (Royseven song), from the album ” The Art of Insincerity, Genre, Rock", "wikidata entities": [ { "id": "Q7375238", "label": "Royseven" }, { "id": "Q188451", "label": "music genre" }, { "id": "Q7714454", "label": "The Art of Insincerity" }, … ], "disambiguations": [ [ "The Art of Insincerity", "Q7714454" ], [ "Royseven", "Q7375238" ], [ "Genre", "Q188451" ], … ], "retrieved for entity": { "id": "Q7085553" }, "source": "info" }
Table 8: Evidence examples with full annotation information.
A.2 CompMix-IR QA Examples
We present some question-answer examples from the CompMix-IR dataset in Table 9, while Table 10 provides a QA example with full annotation information. Table  11 shows the statistics of the CompMix-IR QA set.
@@ 19 August 1988
Question: Who was the voice actor for Meg Griffin in Family Guy? Answer:
Mila Kunis
Annotation information: {"question id": "5136", "question": "Who was the voice actor for Meg Griffin in Family Guy?", "domain": "tvseries", "entities": [{"id": "Q908772", "label": "Meg Griffin"}, {"id": "Q5930", "label": "Family Guy"}], "answers": [{"id": "Q37628", "label": "Mila Kunis"}], "answer text": "Mila Kunis", "answer src": "kg"}
Table 10: A QA example with full annotation information.
@@ Appendix E Retrieving Robustness of UniHGKR
Appendix F Additional Ablation Studies
F.1 Experiments under the Unsupervised Setting
We conduct experiments under the unsupervised setting (i.e., after training in Stage 1 and Stage 2) in retrieval scenario 1, and the results are shown in Table 14. From these results, we can clearly observe the performance gains brought by each stage to the model’s retrieval capabilities. Overall, the alignment training in Stage 2 provides more significant gains compared to the pretraining in Stage 1. After training in Stage 2, the unsupervised model achieves a respectable 73.52 in Hit@100.
Method Hit@5 Hit@10 Hit@100 MRR@100 Bert-base-uncased 6.55 10.93 37.19 5.04 After Stage 1 9.26 14.47 49.78 6.76 Abs. gain +2.71 +3.54 +12.59 +1.72 After Stage 2 16.03 25.54 73.52 12.10 Abs. gain +6.77 +11.07 +23.74 +5.34 After Stage 3* 32.38 45.55 85.75 22.57 Abs. gain +16.35 +20.01 +12.23 +10.47
Table 14: The performance on retrieval scenario 1 after different training stages. Among them, ‘After Stage 1’ and ‘After Stage 2’ can be regarded as the performance in the unsupervised setting. ‘After Stage 3*’ represents our UniHGKR-base model. ‘Abs. gain’ represents the absolute improvement in performance after each training stage.
F.2 The Impact of Instructions for Retrieving from Specific Sources.
[…]
Method Instructions KG-Hit Text-Hit Table-Hit Info-Hit UniHGKR-base $I_{\text{All}}$ 68.60 65.70 28.76 57.34 UniHGKR-base $I_{\tau}$ 75.43 70.30 41.24 66.21 Abs. gain +6.83% +4.60% +12.48% +8.87%
Table 15: Performance of retrieving specific knowledge types with different instructions in retrieval scenario 2. ‘Abs. gain’ refers to the performance improvement brought by using instruction $I_{\tau}$ compared to $I_{\text{All}}$.
Model Size Vector Dim. Avg. Embed Time (100 evd) Avg. Retrieve Time (100 ques) UniHGKR-Base 109M 768 0.46 s 4.35 s UniHGKR-7B 7B 4096 1.54 s 53.29 s
Table 16: Time efficiency comparison between UniHGKR-Base and UniHGKR-7B. The experiment was conducted on a single V100-32G GPU on the CompMix-IR. The data are the average values of three runs of the experiment for 100 pieces of evidence or 100 questions.
F.3 Efficiency of the Proposed Models
For retrieval tasks, efficiency is as important as accuracy. The time cost of retrieval tasks lies in two parts: (1) Embedding, (2) Retrieving. The factor affecting the first part ‘Embedding’ is the parameter scale of the dense embedder. So, the parameter scales of the baselines and UniHGKR models are shown in Table 3 and Table 5. The efficiency of the second part ‘Retrieving’ is affected by the dimension of the vector generated by the retriever. We added an experiment to show the time efficiency difference between the UniHGKR-Base model and the UniHGKR-7B model, as shown in Table 16. The embedding and retrieving average time costs for UniHGKR-7B are 3.35 and 12.25 times longer than those for UniHGKR-Base, respectively. Note that during retrieval, we did not use fast vector retrieval libraries such as Faiss Johnson et al. (2019) but instead performed a naive KNN Steinbach and Tan (2009) computation.
Appendix G Detailed Description of UniHGKR-7B Adaptation
@@
In Stages 1 and 2 of our UniHGKR-7B training, our input sequence $S$ is the linearized structured data $d_{i}$. We adapt EBAE to reconstruct $d_{i}$ and EBAR to predict the corresponding natural language sentence $t_{i}$. Here, $\langle d_{i},t_{i}\rangle$ is from the Data-Text Pairs $\mathcal{D}$. This process essentially implements Stages 1 and 2 of our UniHGKR training framework: establishing an effective representation space for heterogeneous knowledge. For task fine-tuning (Stage 3), we use the same training methods as the UniHGKR-base models (BERT-based), including the instruction set and positive/negative sampling strategies (see Section 4.2).